OpenStudy (anonymous):

A cat needs to figure out if it can jump across a ledge onto another legge. it weights 1 kilo and the ledge is 2 meters away. How much force does he need to make it?

OpenStudy (anonymous):

*ledge

OpenStudy (anonymous):

use projectile equations to find the velocity of jump and then just use newtons second law that states force is nothing but change in momentum...

OpenStudy (anonymous):

a cat dont kno about newton :P

OpenStudy (anonymous):

i thought you were finding the answer... so i just told you the method? are you asking how the cat knows how much force it has to apply ?

OpenStudy (anonymous):

trick question a cat doesn't know about newton or projectiles

OpenStudy (anonymous):

will you please repeat your question ?

OpenStudy (anonymous):

the cat does know about newtons laws!!! sure it may not know it by that name but nonetheless, it knows! Newton's laws are hardwired into it's brain & form a part of its intuition. So.... :P

OpenStudy (anonymous):

By the way, do you want to tricking us?

OpenStudy (anonymous):

As much force as it could take out. :P

OpenStudy (anonymous):

if the jump makes an angle of theta with horizontal range =usquare sine2theta divided by g which is equal to 2 meters in ur question....since we assume cat is stationary before jumping he has to attain an initial velocity of so much....assuming cat knows the maximum range is obtained by jumping at 45 degrees,...he has to accelerate to this velocity before jumping....force=mass*acceleration vf=v0+at here v0=0 aceeleration =vf/t i think i dont know ....just thinking out loud.....someone may give a bette r explanation

OpenStudy (anonymous):

ok force =change in momentum since mass of cat doesnt change mass of cat (vf-v0) mass of cat * initial velocity

OpenStudy (anonymous):

maximum range =(v0)square /g when 45 degrees theta

OpenStudy (anonymous):

maybe i confused u am a lil rusty myself